# Exploring Concept Depth: How Large Language Models Acquire Knowledge and Concepts at Different Layers?

### v5

@@ Abstract
Mingyu Jin1, Qinkai Yu2, Jingyuan Huang1, Qingcheng Zeng3, Zhenting Wang1, Wenyue Hua1, Haiyan Zhao4, Kai Mei1, Yanda Meng5, Kaize Ding3, Fan Yang6, Mengnan Du4, Yongfeng Zhang1 1Rutgers University, 2University of Liverpool, 3Northwestern University, 4New Jersey Institute of Technology, 5University of Exeter, 6Wake Forest University {mingyu.jin, chy.huang, yongfeng.zhang}@rutgers.edu , sgqyu9@liverpool.ac.uk , mengnan.du@njit.edu ,
footnotetext: **Mingyu, Qinkai, Jingyuan, and Qingcheng are the main contributors. The order of the authors is determined by flipping a coin.**
@@ 1 Introduction
LLMs such as GPT-4 Achiam et al. (2023) and LLaMA Touvron et al. (2023) have impressive generation and reasoning capabilities Chang et al. (2023); Su et al. (2024a, c). It is widely accepted that these models embed substantial knowledge in their parameters, with performance improving as the number of parameters increases Ju et al. (2024), also known as emergent abilities Wei et al. (2022a). For instance, GPT-3 Brown et al. (2020) shows a large increase in performance after scaling up to 13B parameters, and a similar phenomenon was also observed for LLaMA Thoppilan et al. (2022) after exceeding 68B parameters Wei et al. (2022b). However, it is not well understood how LLMs accurately grasp the concept of knowledge. In this paper, we investigate the following research question: Can shallow layers in LLMs capture meaningful features of simple knowledge, while complex concepts need deeper layers to capture their meaningful features? like Figure 2. We hope to explore the connection between the depth of language models’ neural networks and their conceptual understanding ability by studying this question.
@@
Our general framework is shown in Figure 1. We trained independent linear probes for each layer of LLMs to predict the binary label, thereby determining the optimal performance achievable with the representations of each respective layer.
Drawing from our empirical findings, we propose the notion of “Concept Depth” as a novel metric to evaluate the capability of different models in comprehending varying levels of knowledge across their layers. This is the first time such a concept has been introduced in the relevant literature.
@@ 2.1 Concepts Representation in DNNs
Identifying similarities across various examples to form concepts, plays a crucial role in human decision-making Armstrong et al. (1983). Many studies have explained DNNs’ (Deep Neural Networks) decision-making based on a conceptual perspective, describing the global behavior of DNNs in a human-understandable way Su et al. (2022); Räz (2023); Ren et al. (2023); Deng et al. (2021). For example, Yeh et al. (2019) demonstrated that DNNs exhibit conceptual representations through the activation patterns observed in their hidden or output layers. Further, Räz (2023) indicates that DNNs learned not only conceptual representations of predicted categories but also indirect concepts that contribute to the prediction. A notable study reveals the existence of a representation bottleneck, highlighting a cognitive disparity between DNNs and humans. This phenomenon is characterized by DNNs’ tendency to grasp representations of concepts that are either too simple or overly complex, while they often struggle with acquiring representations of concepts of moderate complexity Deng et al. (2021). Motivated by previous work, our paper aims to cover concepts of different complexities to understand concepts within LLMs further.
@@ 2.3 Explorations of Interpretability in LLMs
Many related studies have deconstructed the inner layers of LLMs from various perspectives to understand the mechanisms inside such models Zhao et al. (2024); jin2024impactFan et al. (2024) computes stopping signals by evaluating key features to early stop the LLM inference and get the internal performance of the LLMs, concluding that not all layers are necessary. Through pruning the LLMs, Gromov et al. (2024) found that the parameters of some layers were not utilized correctly. Men et al. (2024) also shows a high level of redundancy in the LLMs’ architecture. Probes trained with logistic regression are a well-established method Alain and Bengio (2016) that has been applied in classifying the truthfulness of LLMs and has been validated in many studies Marks and Tegmark (2023); Azaria and Mitchell (2023); Li et al. (2024). The latest work detects different layers in the Llama series responding to facts or counterfactuals by probing techniques Ju et al. (2024). Inspired by these works, we propose Concept Depth to summarize these phenomena. Our work focuses on the Concept Depth that appears in the LLMs, analyzing it experimentally by training linear classifier probes, which makes our work different from others.
@@ 4 Experimental Setting
Our experiments used nine datasets containing three aspects (emotion understanding, reasoning, and fact-checking). We categorized these nine datasets from easy to complex levels according to the performance of LLaMA3-8B-Instruct Dubey et al. (2024) on each dataset (see Section 4.2.1) to anchor the difficulty of the datasets. Specifically, the datasets in which the linear probes can obtain high classification accuracy at the initial or middle depth of the LLMs are categorized as easy levels. Other datasets where linear probes can only accurately classify at a deeper layer of the model or even fail to classify accurately are categorized as complex levels. The overall accuracy of these datasets on the LLaMA3-8B-Instruct model was consistent with the probe results and had a significant correlation. In Section 4.1, we introduce the LLMs used for experiments. The nine datasets are described in Section 4.2.
@@ 4.2.1 Anchoring Difficulties of Each Dataset
To ascertain the learning difficulty of each dataset, we have utilized the LLaMA3-8B-Instruct Dubey et al. (2024) model. Our approach involves testing each sample in the datasets as a binary classification problem via a prompting way. The model generates a response for each sample, from which we infer a judgment, categorizing it as either "Yes" or "No". By comparing these judgments with the actual labels, we compute the accuracy for each dataset.
@@
Table 2: Accuracy on nine datasets based on LLaMA3-8b-Instruct.
@@
Coinflip 0.5080
Counterfact 0.5277
Common 0.5606
Sarcasm 0.6575
StrategyQA 0.7035
@@
Cities 0.7687
STSA 0.9030
IMDb 0.9365

### v6

@@ Abstract
Mingyu Jin1, Qinkai Yu2, Jingyuan Huang1, Qingcheng Zeng3, Zhenting Wang1, Wenyue Hua1, Haiyan Zhao4, Kai Mei1, Yanda Meng2, Kaize Ding3, Fan Yang5, Mengnan Du4, Yongfeng Zhang1 1Rutgers University, 2University of Exeter, 3Northwestern University, 4New Jersey Institute of Technology, 5Wake Forest University, {mingyu.jin, chy.huang, yongfeng.zhang}@rutgers.edu , qy269@exeter.ac.uk , mengnan.du@njit.edu ,
footnotetext: **Mingyu, Qinkai, Jingyuan, and Qingcheng are the main contributors. The order of these authors is determined by flipping a coin.**
@@ 1 Introduction
LLMs such as GPT-4 Achiam et al. (2023) and LLaMA Touvron et al. (2023) have impressive generation and reasoning capabilities Chang et al. (2023); Su et al. (2024a, c). It is widely accepted that these models embed substantial knowledge in their parameters, with performance improving as the number of parameters increases Ju et al. (2024); Jin et al. (2024a); Zhou et al. (2024), also known as emergent abilities Wei et al. (2022a). For instance, GPT-3 Brown et al. (2020) shows a large increase in performance after scaling up to 13B parameters, and a similar phenomenon was also observed for LaMDA Thoppilan et al. (2022) after exceeding 68B parameters Wei et al. (2022b). However, it is not well understood how LLMs accurately grasp the concept of knowledge. In this paper, we investigate the following research question: Can shallow layers in LLMs capture meaningful features of simple knowledge, while complex concepts need deeper layers to capture their meaningful features? like Figure 2. We hope to explore the connection between the depth of language models’ neural networks and their conceptual understanding ability by studying this question.
@@
Our general framework is shown in Figure 1. We trained independent linear probes for each layer of LLMs to predict the binary label, thereby determining the optimal performance achievable with the representations of each respective layer. Drawing from our empirical findings, we propose the notion of “Concept Depth” as a novel metric to evaluate the capability of different models in comprehending varying levels of knowledge across their layers. This is the first time such a concept has been introduced in the relevant literature.
@@ 2.1 Concepts Representation in DNNs
Identifying similarities across various examples to form concepts, plays a crucial role in human decision-making Armstrong et al. (1983). Many studies have explained DNNs’ (Deep Neural Networks) decision-making based on a conceptual perspective, describing the global behavior of DNNs in a human-understandable way Su et al. (2022); Räz (2023); Ren et al. (2023); Deng et al. (2021); Wen et al. (2024a); Wen (2024); Wen et al. (2024b). For example, Yeh et al. (2019) demonstrated that DNNs exhibit conceptual representations through the activation patterns observed in their hidden or output layers. Further, Räz (2023) indicates that DNNs learned not only conceptual representations of predicted categories but also indirect concepts that contribute to the prediction. A notable study reveals the existence of a representation bottleneck, highlighting a cognitive disparity between DNNs and humans. This phenomenon is characterized by DNNs’ tendency to grasp representations of concepts that are either too simple or overly complex, while they often struggle with acquiring representations of concepts of moderate complexity Deng et al. (2021). Motivated by previous work, our paper aims to cover concepts of different complexities to understand concepts within LLMs further.
@@ 2.3 Explorations of Interpretability in LLMs
Many related studies have deconstructed the inner layers of LLMs from various perspectives to understand the mechanisms inside such models Zhao et al. (2024); Jin et al. (2024b)Fan et al. (2024) computes stopping signals by evaluating key features to early stop the LLM inference and get the internal performance of the LLMs, concluding that not all layers are necessary. Through pruning the LLMs, Gromov et al. (2024) found that the parameters of some layers were not utilized correctly. Men et al. (2024) also shows a high level of redundancy in the LLMs’ architecture. Probes trained with logistic regression are a well-established method Alain and Bengio (2016) that has been applied in classifying the truthfulness of LLMs and has been validated in many studies Marks and Tegmark (2023); Azaria and Mitchell (2023); Li et al. (2024). The latest work detects different layers in the Llama series responding to facts or counterfactuals by probing techniques Ju et al. (2024). Inspired by these works, we propose Concept Depth to summarize these phenomena. Our work focuses on the Concept Depth that appears in the LLMs, analyzing it experimentally by training linear classifier probes, which makes our work different from others.
@@ 4 Experimental Setting
Our experiments used nine datasets containing three aspects (emotion understanding, reasoning, and fact-checking). We categorized these nine datasets from easy to complex levels according to the performance of LLaMA3-8B-Instruct Dubey et al. (2024), GPT-4o-mini OpenAI (2024), and QWen2-7B-Instruct Yang et al. (2024) on each dataset (see Section 4.2.1) to anchor the difficulty of the datasets. Specifically, the datasets in which the linear probes can obtain high classification accuracy at the initial or middle depth of the LLMs are categorized as easy levels. Other datasets where linear probes can only accurately classify at a deeper layer of the model or even fail to classify accurately are categorized as complex levels. The average accuracy of these datasets on the three models was consistent with the probe results and had a significant correlation. In Section 4.1, we introduce the LLMs used for experiments. The nine datasets are described in Section 4.2.
@@ 4.2.1 Anchoring Difficulties of Each Dataset
To ascertain the learning difficulty of each dataset, we have utilized the LLaMA3-8B-Instruct Dubey et al. (2024), GPT-4o-mini OpenAI (2024), and QWen2-7B-Instruct model Yang et al. (2024). Our approach involves testing each sample in the datasets as a binary classification problem via a prompting way. The model generates a response for each sample, from which we infer a judgment, categorizing it as either "Yes" or "No". By comparing these judgments with the actual labels, we compute the accuracy for each dataset.
@@
Table 2: Average accuracy on nine datasets based on LLaMA3-8b-Instruct, GPT-4o-mini and QWen2-7B-Instruct. Accuracy based on each model is shown in Table 5.
@@
Coinflip 0.5920
Common 0.6487
Sarcasm 0.6597
StrategyQA 0.6969
Counterfact 0.7126
@@
STSA 0.9116
Cities 0.9204
IMDb 0.9380
@@ 7 Limitations
8 Acknowledgement
We thank Taowen Wang, Fei Sun, Wujiang Xu, and Guangyan Sun for their valuable discussions and suggestions during the project.
@@ A.4 Examples of Each Dataset
Table 5: Accuracy on nine datasets based on LLaMA3-8b-Instruct, GPT-4o-mini and QWen2-7B-Instruct.
Dataset LLaMA3-8B-Instruct GPT-4o-mini QWen2-7B-Instruct Average
Coinflip 0.5080 0.7620 0.5060 0.5920
Common 0.5606 0.6905 0.6950 0.6487
Sarcasm 0.6575 0.6770 0.6445 0.6597
StrategyQA 0.7035 0.8803 0.5069 0.6969
Counterfact 0.5277 0.7990 0.8110 0.7126
Hateeval 0.7640 0.7300 0.7952 0.7631
STSA 0.9030 0.9211 0.9108 0.9116
Cities 0.7687 0.9973 0.9953 0.9204
IMDb 0.9365 0.9370 0.9405 0.9380